OpenStudy (goformit100):

If p is prime, then (p-1)! + 1 =0 mod p Prove it

OpenStudy (anonymous):

Wow interestingly all your questions are the ones I have only just studied a few months earlier! :) Here you go: Wilson's Theorem! I suggest you read up on this theorem. I think this is a good proof: (taken from http://primes.utm.edu/notes/proofs/Wilsons.html) It is easy to check the result when p is 2 or 3, so let us assume p > 3. If p is composite, then its positive divisors are among the integers 1, 2, 3, 4, ... , p-1 and it is clear that gcd((p-1)!,p) > 1, so we can not have (p-1)! = -1 (mod p). However if p is prime, then each of the above integers are relatively prime to p. So for each of these integers a there is another b such that ab = 1 (mod p). It is important to note that this b is unique modulo p, and that since p is prime, a = b if and only if a is 1 or p-1. Now if we omit 1 and p-1, then the others can be grouped into pairs whose product is one showing 2*3*4.....(p-2) = 1 (mod p) (or more simply (p-2)! = 1 (mod p)). Finally, multiply this equality by p-1 to complete the proof. There you go! Hope that helped. Have a great day!

OpenStudy (goformit100):

Thankyou